A manager wishes to build a 3-sigma range chart for a process. The sample size is five, the mean of sample means is 28.5, and the average range is 3.9. From Table S6.1, the appropriate value of D3 is 0, and D4 is 2.115. What are the UCL and LCL, respectively, for this range chart?
1
Expert's answer
2020-12-04T07:46:57-0500

Upper and lower limits for the Range chart:

"UCLR = D4 \times Rbar =2.115 \times 3.9 = 8.2485 \\

LCLR = D3 \times Rbar =0 \times 3.9=0"

Answer: 8.25 and 0